# BASE-Q: Bias and Asymmetric Scaling Enhanced Rotational Quantization for Large Language Models

### v1

@@ Abstract
Rotations have become essential to state-of-the-art quantization pipelines for large language models (LLMs) by effectively smoothing outliers in weights and activations. However, further optimizing the rotation parameters offers only limited performance gains and introduces significant training overhead: due to rotation parameter sharing, full-model must be loaded simultaneously to enable backpropagation, resulting in substantial memory consumption and limited practical utility. In this work, we identify two fundamental limitations of current rotational quantization methods: (i) rotation fails to align channel means, resulting in wider quantization bounds and increased rounding errors; and (ii) rotation makes the activation distribution more Gaussian-like, increasing energy loss caused by clipping errors. To address these issues, we introduce BASE-Q, a simple yet powerful approach that combines bias correction and asymmetric scaling to effectively reduce rounding and clipping errors. Furthermore, BASE-Q enables blockwise optimization, eliminating the need for memory-intensive full-model backpropagation. Extensive experiments on various LLMs and benchmarks demonstrate the effectiveness of BASE-Q, narrowing the accuracy gap to full-precision models by 50.5%, 42.9%, and 29.2% compared to QuaRot, SpinQuant, and OSTQuant, respectively. The code will be released soon.
@@ 3 Error analysis of existing methods
In this section, we theoretically analyze the errors derived by rotation-based quantization, specifically focusing on the Hadamard transformation. We begin by investigating the error reduction mechanism inherent in this transformation, deriving quantitative expressions for both rounding and clipping errors in Section 3.1. Subsequently, in Sections 3.2 and 3.3, we provide an in-depth analysis of these error sources, uncovering potential components that are amenable to further optimization and providing key insights for methodological improvements. Finally, drawing inspiration from OSTQuantHu2025OstQuantRL , which demonstrates complementary effects of scaling and rotation, we interpret and substantiate this synergy through our error decomposition framework in Section 3.4.
@@ 3.1 Quantization error analysis for Hadamard rotations
where $\boldsymbol{\mathit{g}}\sim\mathcal{N}(\mu,\delta^{2}\mathbf{I})$ represents the “main mass” as a Gaussian component, and each outlier channel $i\in\mathcal{O}$ is modeled as a scaled one-hot vector of amplitude $a_{i}\delta$ with $a_{i}\gg 1$ and $|\mathcal{O}|\ll n$. When an orthogonal Hadamard rotation $\boldsymbol{\mathit{H}}$ is applied (where each entry is ±$\pm n^{-\frac{1}{2}}$), the transformed activation becomes:
@@
[…]
@@ 3.2 Rounding errors from misaligned channel means
Figure 1: (a) Raw activation distribution of the first MLP block in Llama3-8B. (b) Hadamard rotation suppresses outliers but leaves residual inter-channel misalignment. (c) Learned rotations also fail to address this misalignment. (d) Our bias correction eliminates most inter-channel mean variability, thereby reducing $Var(\mu_{j})$, which is the main rounding error component as formalized in Equations 5 and 6. More visualizations are in Appendix B.
@@
where $\sigma^{2}_{j}$ and $\mu_{j}$ are the variance and mean of the $j$-th channel, respectively. The first term, the average channel variance, is invariant under orthogonal transformations and dictated by the data distribution. In contrast, the second term, $Var(\mu_{j})$ —the variance of channel means—can remain substantial in LLMs that use RMSNorm, since this normalization does not enforce alignment across channel means. As shown in Figure 1, hadamard rotations are effective at attenuating outliers but leave the means of individual channels misaligned. Surprisingly, according to Equation 6, we found that $Var(\mu_{j})$ causes up to 85% layer rounding error in Qwen2.5-3B (as shown in Figure 2; additional details for more models are provided in Appendix C).
@@ 3.3 Clipping eliminates non-negligible energy
[…]
@@
[…]
where $P(x)$ is the probability density function. 18.4% of the energy is non-negligible in the context of deep neural networks. When this clipping is applied layer by layer throughout a deep network, the accumulated energy loss becomes substantial. The empirical consequence of this breakdown is illustrated in Figure 3, where the distribution of the layer output significantly mismatches the original. Moreover, in practical LLM architectures, using globally or per-layer optimized clipping bounds cannot fully resolve this energy loss, as joint optimization across all layers is intractable. This limitation explains the observed diminishing returns of aggressive clipping in rotation-based quantization schemes, even with parameter tuning.
@@ 4 BASE-Q
As discussed in Sections 3.2 and 3.3, while rotation-based quantization methods effectively smooth outliers, they exhibit three notable limitations: (a) Channel-mean variance cannot be eliminated via rotation-only optimization, leading to a persistent rounding error term (as formalized in Equations 5 and 6). (b) Implementing flexible and optimal activation clipping schemes within standard rotation-based PTQ is challenging, resulting in cumulative energy loss across layers. (c) Optimizing full-network rotation parameters requires prohibitive GPU memory and computational resources, especially for large models. To address these issues, we propose BASE-Q (Bias and Asymmetric Scaling Enhanced Quantization), a lightweight yet powerful quantization framework that strengthens rotation-based methods through two key innovations: explicit bias correction to fully remove channel mean variance, and asymmetric scaling to improve clipping and quantizer fit at each block. The blockwise optimization capability avoids the heavy cost and complexity of full-model rotation learning, thereby enabling efficiency and enhanced overall quantization performance.
@@
Figure 4 provides a schematic of our BASE-Q. The core optimized transformations within each transformer block include the global shared residual rotation ($\boldsymbol{\mathit{R}}_{res}$), the blockwise value rotation ($\boldsymbol{\mathit{R}}_{v}$), and the online rotations ($\boldsymbol{\mathit{R}}_{qk}$ and $\boldsymbol{\mathit{R}}_{down}$). $\boldsymbol{\mathit{R}}_{res}$ is applied on the residual pathway, which can be fused into linear weights without extra inference overhead. Because the residual path connects all blocks, a single shared rotation is used throughout. As our method further removes channel mean variance and optimizes quantization for activations, it is preferable to set $\boldsymbol{\mathit{R}}_{res}$ to primarily optimize weight statistics. Following Section 3.1, we compute the PCA transform $\boldsymbol{\mathit{U}}$ across all related linear weights, and compose it with the Hadamard transformation:
@@
where $\boldsymbol{\mathit{WR}}$ denotes rotated weights and $b$ is the usual layer bias. The learnable bias correction terms—including $B^{c}_{qkv}$, $B^{c}_{o}$, $B^{c}_{up}$ and $B^{c}_{down}$, which incur negligible parameter overhead (less than 0.1% of original model size), and are highly efficient to optimize.
@@
For the O and Down linear layers, we introduce both global symmetric $\boldsymbol{\mathit{S}}$ and per-quantizer asymmetric scaling $\boldsymbol{\mathit{S}}^{a}$. Asymmetric scaling $S^{a}$ can adjusts the quantization range at inference time to better fit activation statistics and clipping needs:
@@
Combining these techniques, our blockwise optimization strategy jointly optimizes $\boldsymbol{\mathit{R}}_{v}$, $\boldsymbol{\mathit{B}}^{c}_{i}$,$\boldsymbol{\mathit{S}}_{j}$, $\boldsymbol{\mathit{S}}^{a}_{i}$, and activation clipping threshold $\alpha_{i}$ per block via minimizing the MSE between the floating-point and quantized outputs:
@@ 5.1 Main results
Section 5.1 presents the overall performance of BASE-Q and baseline methods under W4A4KV4 quantization, reporting perplexity on Wikitext-2 and the average accuracy across nine zero-shot tasks. Since BASE-Q is explicitly designed to tackle the most challenging 4-bit activation quantization, we focus solely on the W4A4KV4 configuration. Across all evaluated models, BASE-Q consistently achieves superior results, with an average accuracy drop of only 3.40% from full precision. In comparison, the average accuracy drops for QuaRot, SpinQuant, and OSTQuant are 6.87%, 5.95%, and 4.80%, respectively (averaged over eight supported models). This translates to BASE-Q narrowing the gap to full-precision by 50.5%, 42.9%, and 29.2% relative to QuaRot, SpinQuant, and OSTQuant, respectively. Notably, BASE-Q performs exceptionally well on Qwen2.5-3B, where all existing methods suffer severe degradation (perplexity increases by more than 100%). These results strongly validate our theoretical insights and demonstrate the practical effectiveness of our proposed bias correction and asymmetric scaling strategies.
Table 1: Comparison on Wikitext-2 perplexity and 9 zero-shot benchmark tasks. All baseline results—QuaRot, SpinQuant, and OSTQuant—are reproduced using their official open-source implementations, with necessary modifications to support the Qwen model, which differs from LLaMA mainly by including attention bias. As the official OSTQuant repository does not support FSDP (Fully Sharded Data Parallel) training, its results are limited to models with up to 14B parameters. Complete results are provided in Appendix A.
@@
[…]
@@
[…]
@@ 5.2 Prefill acceleration on GPU
To evaluate real-world acceleration, we benchmark INT4 quantization during the compute-bound prefill stage on an NVIDIA 3090 GPU. All experiments use a sequence length of 2048, with batch sizes varying from 1 to 64. 4-bit matrix multiplications are implemented by the NVIDIA’s Cutlass library, while other custom operators are written in Triton for flexibility and speed. To minimize overhead, we fuse the online bias and scaling introduced by BASE-Q within quantization and de-quantization into a single Triton kernel (Figure 32), rendering the additional costs of bias and scaling computation negligible. For the online Hadamard transform, we find that the fast algorithm fast-hadamard-transform adopted in Quarot, which uses the recursive Cooley-Tukey method to achieve $O(nlogn)$ complexity, does not fully leverage tensor core parallelism on modern GPUs. To address this, we implement the Hadamard transform as a larger matrix multiplication in Triton, better utilizing available compute. For query and key projections (with small dimensions, e.g., 128×128), we perform the transformation directly as a matrix multiplication. For the much larger down layer, we only apply one single Cooley-Tukey step, ensuring all compute-intensive kernels operate on manageable, high-parallelism sub-matrices. As shown in Figure 31, our approach yields 2.1×2.1$\times$2.1 × to 2.4×2.4$\times$2.4 × acceleration across all batch sizes compared with standard FP16. Notably, our optimizations incur only minimal overhead versus pure INT4 quantization without any online operations.
@@ 5.3 Ablation study
We perform a systematic ablation study on Qwen2.5-3B, Llama2-7B, and Llama3-8B to assess the effects of different quantization strategies in BASE-Q, as well as those from Quarot, SpinQuant, and OSTQuant. Our experiments as shown in Section 5.3 indicate that bias correction yields substantial perplexity reductions for Qwen2.5-3B and Llama3-8B, highlighting the importance of addressing inter-channel bias in these models. In contrast, its effect on Llama2-7B is marginal. Notably, asymmetric scaling delivers consistent improvements across all three models, confirming its effectiveness in mitigating the adverse impact of clipping.
@@
Method Fixed Rotation Learned Rotation Bias Corect. Unpaired Scale Scale Qwen2.5-3B Llama-2-7B Llama-3-8B
@@
[…]
@@
[…]

### v2

@@ Abstract
Rotations have become essential to state-of-the-art quantization pipelines for large language models (LLMs) by effectively smoothing outliers in weights and activations. However, further optimizing the rotation parameters offers only limited performance gains and introduces significant training overhead: due to rotation parameter sharing, full-model must be loaded simultaneously to enable backpropagation, resulting in substantial memory consumption and limited practical utility. In this work, we identify two fundamental limitations of current rotational quantization methods: (i) rotation fails to align channel means, resulting in wider quantization bounds and increased rounding errors; and (ii) rotation makes the activation distribution more Gaussian-like, increasing energy loss caused by clipping errors. To address these issues, we introduce BASE-Q, a simple yet powerful approach that combines bias correction and asymmetric scaling to effectively reduce rounding and clipping errors. Furthermore, BASE-Q enables blockwise optimization, eliminating the need for memory-intensive full-model backpropagation. Extensive experiments on various LLMs and benchmarks demonstrate the effectiveness of BASE-Q, narrowing the accuracy gap to full-precision models by 50.5%, 42.9%, and 29.2% compared to QuaRot, SpinQuant, and OSTQuant, respectively. The code can be found at https://github.com/Heliulu/BASE-Q.
@@ 3 Error analysis of existing methods
In this section, we theoretically analyze the errors derived by rotation-based quantization, specifically focusing on the Hadamard transformation. We begin by investigating the error reduction mechanism inherent in this transformation, deriving quantitative expressions for both rounding and clipping errors in Section˜3.1. Subsequently, in Sections˜3.2 and 3.3, we provide an in-depth analysis of these error sources, uncovering potential components that are amenable to further optimization and providing key insights for methodological improvements. Finally, drawing inspiration from OSTQuantHu2025OstQuantRL , which demonstrates complementary effects of scaling and rotation, we interpret and substantiate this synergy through our error decomposition framework in Section˜3.4.
@@ 3.1 Quantization error analysis for Hadamard rotations
where $\boldsymbol{\mathit{g}}\sim\mathcal{N}(\mu,\delta^{2}\mathbf{I})$ represents the “main mass” as a Gaussian component, and each outlier channel $i\in\mathcal{O}$ is modeled as a scaled one-hot vector of amplitude $a_{i}\delta$ with $a_{i}\gg 1$ and $|\mathcal{O}|\ll n$. When an orthogonal Hadamard rotation $\boldsymbol{\mathit{H}}$ is applied (where each entry is $\pm n^{-\frac{1}{2}}$), the transformed activation becomes:
@@
  • Clipping error, defined by the mass outside quantization bounds: $\mathbb{E}[\|\varepsilon_{\text{clip}}\|_{2}^{2}]=\int^{z}_{-\infty}x^{2}P(x)dx+\int^{+\infty}_{z+\Delta}x^{2}P(x)dx$, where $\Delta=s(2^{b}-1)$ and $P(x)$ is the empirical channel distribution.
@@ 3.2 Rounding errors from misaligned channel means
Figure 1: (a) Raw activation distribution of the first MLP block in Llama3-8B. (b) Hadamard rotation suppresses outliers but leaves residual inter-channel misalignment. (c) Learned rotations also fail to address this misalignment. (d) Our bias correction eliminates most inter-channel mean variability, thereby reducing $Var(\mu_{j})$, which is the main rounding error component as formalized in Equations˜5 and 6. More visualizations are in Appendix˜B.
@@
where $\sigma^{2}_{j}$ and $\mu_{j}$ are the variance and mean of the $j$-th channel, respectively. The first term, the average channel variance, is invariant under orthogonal transformations and dictated by the data distribution. In contrast, the second term, $Var(\mu_{j})$ —the variance of channel means—can remain substantial in LLMs that use RMSNorm, since this normalization does not enforce alignment across channel means. As shown in Figure˜1, hadamard rotations are effective at attenuating outliers but leave the means of individual channels misaligned. Surprisingly, according to Equation˜6, we found that $Var(\mu_{j})$ causes up to 85% layer rounding error in Qwen2.5-3B (as shown in Figure˜2; additional details for more models are provided in Appendix˜C).
@@ 3.3 Clipping eliminates non-negligible energy
$\theta^{*}=\underset{\theta}{\arg\min}\ \mathbb{E}[\|\boldsymbol{\mathit{x}}-\theta-F_{clip}(F_{round}(\frac{\boldsymbol{\mathit{x}}-\theta}{2\theta/(2^{b}-1)}),0,2^{b}-1)\|_{2}]\approx 2.2\sigma$ (7)
@@
$\mathbb{E}[\|\varepsilon_{\text{clip}}\|_{2}]=\int^{-2.2\sigma}_{-\infty}x^{2}P(x)dx+\int^{+\infty}_{+2.2\sigma}x^{2}P(x)dx\approx 18.4\%,$ (8)
where $P(x)$ is the probability density function. 18.4% of the energy is non-negligible in the context of deep neural networks. When this clipping is applied layer by layer throughout a deep network, the accumulated energy loss becomes substantial. The empirical consequence of this breakdown is illustrated in Figure˜3, where the distribution of the layer output significantly mismatches the original. Moreover, in practical LLM architectures, using globally or per-layer optimized clipping bounds cannot fully resolve this energy loss, as joint optimization across all layers is intractable. This limitation explains the observed diminishing returns of aggressive clipping in rotation-based quantization schemes, even with parameter tuning.
@@ 4 BASE-Q
As discussed in Sections˜3.2 and 3.3, while rotation-based quantization methods effectively smooth outliers, they exhibit three notable limitations: (a) Channel-mean variance cannot be eliminated via rotation-only optimization, leading to a persistent rounding error term (as formalized in Equations˜5 and 6). (b) Implementing flexible and optimal activation clipping schemes within standard rotation-based PTQ is challenging, resulting in cumulative energy loss across layers. (c) Optimizing full-network rotation parameters requires prohibitive GPU memory and computational resources, especially for large models. To address these issues, we propose BASE-Q (Bias and Asymmetric Scaling Enhanced Quantization), a lightweight yet powerful quantization framework that strengthens rotation-based methods through two key innovations: explicit bias correction to fully remove channel mean variance, and asymmetric scaling to improve clipping and quantizer fit at each block. The blockwise optimization capability avoids the heavy cost and complexity of full-model rotation learning, thereby enabling efficiency and enhanced overall quantization performance.
@@
Figure˜4 provides a schematic of our BASE-Q. The core optimized transformations within each transformer block include the global shared residual rotation ($\boldsymbol{\mathit{R}}_{res}$), the blockwise value rotation ($\boldsymbol{\mathit{R}}_{v}$), and the online rotations ($\boldsymbol{\mathit{R}}_{qk}$ and $\boldsymbol{\mathit{R}}_{down}$). $\boldsymbol{\mathit{R}}_{res}$ is applied on the residual pathway, which can be fused into linear weights without extra inference overhead. Because the residual path connects all blocks, a single shared rotation is used throughout. As our method further removes channel mean variance and optimizes quantization for activations, it is preferable to set $\boldsymbol{\mathit{R}}_{res}$ to primarily optimize weight statistics. Following Section˜3.1, we compute the PCA transform $\boldsymbol{\mathit{U}}$ across all related linear weights, and compose it with the Hadamard transformation:
@@
where $\boldsymbol{\mathit{WR}}$ denotes rotated weights and $b$ is the usual layer bias. The learnable bias correction terms—including $\boldsymbol{\mathit{b}}^{c}_{qkv}$, $\boldsymbol{\mathit{b}}^{c}_{o}$, $\boldsymbol{\mathit{b}}^{c}_{up}$ and $\boldsymbol{\mathit{b}}^{c}_{down}$, which incur negligible parameter overhead (less than 0.1% of original model size), and are highly efficient to optimize.
@@
For the O and Down linear layers, we introduce both global symmetric $\boldsymbol{\mathit{s}}$ and per-quantizer asymmetric scaling $\boldsymbol{\mathit{s}}^{a}$. Asymmetric scaling $\boldsymbol{\mathit{s}}^{a}$ can adjusts the quantization range at inference time to better fit activation statistics and clipping needs:
@@
Combining these techniques, our blockwise optimization strategy jointly optimizes $\boldsymbol{\mathit{R}}_{v}$, $\boldsymbol{\mathit{b}}^{c}_{i}$$\boldsymbol{\mathit{s}}_{j}$, $\boldsymbol{\mathit{s}}^{a}_{i}$, and activation clipping threshold $\alpha_{i}$ per block via minimizing the MSE between the floating-point and quantized outputs:
@@ 5.1 Main results
Table˜1 presents the overall performance of BASE-Q and baseline methods under W4A4KV4 quantization, reporting perplexity on Wikitext-2 and the average accuracy across nine zero-shot tasks. Since BASE-Q is explicitly designed to tackle the most challenging 4-bit activation quantization, we focus solely on the W4A4KV4 configuration. Across all evaluated models, BASE-Q consistently achieves superior results, with an average accuracy drop of only 3.40% from full precision. In comparison, the average accuracy drops for QuaRot, SpinQuant, and OSTQuant are 6.87%, 5.95%, and 4.80%, respectively (averaged over eight supported models). This translates to BASE-Q narrowing the gap to full-precision by 50.5%, 42.9%, and 29.2% relative to QuaRot, SpinQuant, and OSTQuant, respectively. Notably, BASE-Q performs exceptionally well on Qwen2.5-3B, where all existing methods suffer severe degradation (perplexity increases by more than 100%). These results strongly validate our theoretical insights and demonstrate the practical effectiveness of our proposed bias correction and asymmetric scaling strategies.
Table 1: Comparison on Wikitext-2 perplexity and 9 zero-shot benchmark tasks. All baseline results—QuaRot, SpinQuant, and OSTQuant—are reproduced using their official open-source implementations, with necessary modifications to support the Qwen model, which differs from LLaMA mainly by including attention bias. As the official OSTQuant repository does not support FSDP (Fully Sharded Data Parallel) training, its results are limited to models with up to 14B parameters. Complete results are provided in Appendix˜A.
@@
W4A4KV4 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki
@@
W4A4KV4 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki
@@ 5.2 Prefill acceleration on GPU
To evaluate real-world acceleration, we benchmark INT4 quantization during the compute-bound prefill stage on an NVIDIA 3090 GPU. All experiments use a sequence length of 2048, with batch sizes varying from 1 to 64. 4-bit matrix multiplications are implemented by the NVIDIA’s Cutlass library, while other custom operators are written in Triton for flexibility and speed. To minimize overhead, we fuse the online bias and scaling introduced by BASE-Q within quantization and de-quantization into a single Triton kernel (Figure˜6), rendering the additional costs of bias and scaling computation negligible. For the online Hadamard transform, we find that the fast algorithm fast-hadamard-transform adopted in Quarot, which uses the recursive Cooley-Tukey method to achieve $O(nlogn)$ complexity, does not fully leverage tensor core parallelism on modern GPUs. To address this, we implement the Hadamard transform as a larger matrix multiplication in Triton, better utilizing available compute. For query and key projections (with small dimensions, e.g., 128×128), we perform the transformation directly as a matrix multiplication. For the much larger down layer, we only apply one single Cooley-Tukey step, ensuring all compute-intensive kernels operate on manageable, high-parallelism sub-matrices. As shown in Figure˜5, our approach yields $2.1\times$ to $2.4\times$ acceleration across all batch sizes compared with standard FP16. Notably, our optimizations incur only minimal overhead versus pure INT4 quantization without any online operations.
@@ 5.3 Ablation study
We perform a systematic ablation study on Qwen2.5-3B, Llama2-7B, and Llama3-8B to assess the effects of different quantization strategies in BASE-Q, as well as those from Quarot, SpinQuant, and OSTQuant. Our experiments as shown in Table˜2 indicate that bias correction yields substantial perplexity reductions for Qwen2.5-3B and Llama3-8B, highlighting the importance of addressing inter-channel bias in these models. In contrast, its effect on Llama2-7B is marginal. Notably, asymmetric scaling delivers consistent improvements across all three models, confirming its effectiveness in mitigating the adverse impact of clipping.
@@
Method
Fixed
Rotation
Learned
Rotation
Bias
Corect.
Unpaired
Scale
Scale Qwen2.5-3B Llama-2-7B Llama-3-8B
@@
${\color[rgb]{.75,.5,.25}\boldsymbol{\mathit{R}}_{res}}\;\boldsymbol{\mathit{R}}_{qk}\;\boldsymbol{\mathit{R}}_{down}$ $\boldsymbol{\mathit{R}}_{v}$ 54.38 -5.95 6.18 +0.06 7.63 -0.19
@@
SpinQuant $\boldsymbol{\mathit{R}}_{qk}\;\boldsymbol{\mathit{R}}_{down}$ ${\color[rgb]{.75,.5,.25}\boldsymbol{\mathit{R}}_{res}}\;\boldsymbol{\mathit{R}}_{v}$ 46.35 5.99 7.49
@@
We further performed ablation studies where, within the BASE-Q framework, global rotations were respectively implemented as a standard Hadamard matrix, a random Hadamard matrix, and a learnable rotation (following SpinQuant’s official codebase). In SpinQuant (cf. Fig. 4 in Liu2024SpinQuantLQ ), the authors observed that using random Hadamard rotations could cause large fluctuations and subpar quantization performance. Based on this observation, SpinQuant advocates for learnable global rotations to stabilize and improve quantization results under their framework.
However, in our ablation studies Table˜3 we found that, across five benchmark LLMs, the quantization metrics showed negligible differences across different rotation types within our BASE-Q framework. This suggests that any potential improvements from learnable global rotations are effectively subsumed by BASE-Q’s bias correction component. Therefore, the additional memory and computational cost introduced by learnable global rotations is unnecessary for maintaining quantization quality in our method.
Table 3: Ablation study on $R_{res}$ choice.
Qwen-2.5 3B Qwen-2.5 14B LLaMA-2 7B LLaMA-3 8B LLaMA-2 13B
W4A4KV4 $R_{res}$ Setting 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki 0-shot9 Wiki
Avg.($\uparrow$) ($\downarrow$) Avg.($\uparrow$) ($\downarrow$) Avg.($\uparrow$) ($\downarrow$) Avg.($\uparrow$) ($\downarrow$) Avg.($\uparrow$) ($\downarrow$)
Full-Precision 64.17 8.03 70.95 5.29 65.22 5.47 68.11 6.14 67.62 4.88
QuaRot Random Hadamard 44.30 69.33 67.23 6.77 61.59 6.12 62.89 7.82 65.03 5.39
SpinQuant Learned Rotation 46.86 46.35 67.29 6.55 61.38 5.99 63.80 7.49 65.63 5.30
BASE-Q Standard Hadamard 58.93 10.43 69.17 6.28 62.08 5.85 65.33 7.13 65.89 5.20
Random Hadamard 58.60 10.43 69.22 6.28 62.24 5.87 65.21 7.14 65.46 5.20
Learned Hadamard 57.95 10.44 68.43 6.26 62.68 5.86 65.01 7.13 65.60 5.20